OpenStudy (anonymous):

Which of the following is equivalent to log32 4? (the 32 is small) pls help, taking a test now

OpenStudy (anonymous):

answers are : 0.4 0.424 0.125 2.5

OpenStudy (aum):

Use the change of base formula. Both 4 and 32 are powers of 2. Use 2 as the new base for the log.

OpenStudy (jdoe0001):

" taking a test now" ? hmmm

OpenStudy (anonymous):

i dont know how to do logs at all... i never learned how to do the change of base formula @aum

OpenStudy (aum):

\( \Huge \log_BA = \frac{\log_cA}{\log_cB} \)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \log_{32}4 =\log_{32}2^2 =2\log_{32}2=2\log_{32}32^{1/5}=}\) \(\normalsize\color{blue}{ 2\log_{32}32^{1/5}= (1/5)(2)\log_{32}32=(2/5)\log_{32}32}\) ••••

OpenStudy (solomonzelman):

This is not the final answer

OpenStudy (solomonzelman):

You got to know how to reduce it. If you can't, then I won't reduce it for you, since you are talking a test. I am not violating more than what I already did.

OpenStudy (aum):

\( \Large \log_{32}4 = \frac{\log_{2}4}{\log_{2}32 } = \frac{?}{?}\)

OpenStudy (anonymous):

1/8 ?

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

oh.. thanks. :)

OpenStudy (aum):

Are you taking a test?

OpenStudy (anonymous):

yes @aum

OpenStudy (imstuck):

we are not allowed to help on tests.